OpenStudy (anonymous):

The rectangular coordinates of the pint (5, pi/3) are... ??? HELP

OpenStudy (ipwnbunnies):

To go to rectangular coordinates, (X,Y): \[X = r*\cos \theta\] \[Y = r* \sin \theta\] Where (r,theta) is the polar point